### v1

# Large Language Model Guided Self-Debugging Code Generation

Abstract.
Automated code generation is gaining significant importance in intelligent computer programming and system deployment. However, current approaches often face challenges in computational efficiency and lack robust mechanisms for code parsing and error correction. In this work, we propose a novel framework, PyCapsule, with a simple yet effective two-agent pipeline and efficient self-debugging modules for Python code generation. PyCapsule features sophisticated prompt inference, iterative error handling, and case testing, ensuring high generation stability, safety, and correctness. Empirically, PyCapsule achieves up to 5.7% improvement of success rate on HumanEval, 10.3% on HumanEval-ET, and 24.4% on BigCodeBench compared to the state-of-art methods. We also observe a decrease in normalized success rate given more self-debugging attempts, potentially affected by limited and noisy error feedback in retention. PyCapsule demonstrates broader impacts on advancing lightweight and efficient code generation for artificial intelligence systems.
Python Code Generation, Self-debugging, Large Language Model, Programming Agent, Artificial Intelligence Automation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Software and its engineering Automatic programmingccs: Computing methodologies Natural language processing
M. Adnan, Z. Xu, and C. C.N. Kuhn
1. Introduction
The evolution of generative artificial intelligence has profoundly influenced multiple domains, notably in automated code generation. Particularly, Large Language Models (LLMs) have demonstrated remarkable proficiency in code generation given natural language prompts, streamlining software development processes and enhancing productivity  (Rasnayaka et al., 2024; Liu et al., 2024; Chowdhery et al., 2024). This transformative potential has sparked significant research in this field. However, several critical challenges emerge as these systems scale to handle increasingly complex programming tasks. Ensuring reliability, accuracy, and autonomy in code generation remains critical but challenging, particularly in error detection and debugging, that align with human developers’ intentions  (Fan et al., 2023; Dou et al., 2024; Lian et al., 2024). Recent research has introduced frameworks consisting of multiple LLM-based agents to emulate human problem-solving capability in code generation. For instance, MapCoder  (Islam et al., 2024), AgentCoder  (Huang et al., 2024b), and LDB  (Zhong et al., 2024) employ agents dedicated to retrieval, planning, coding, and debugging to enhance code quality. While these approaches have shown remarkable improvements, they often require substantial computational resources and complex agent coordination. This poses challenges for real-world deployment and scalability.
To address these concerns, we propose a novel framework, PyCapsule, that enhances the efficiency and accuracy of LLM-based Python code generation through a two-agent architecture. This simple yet effective two-agent interaction ensures functionality independence without overcrowding with AI agents, thereby minimizing its computational consumption. PyCapsule employs a programmer agent responsible for code generation and debugging, interacting with an executor agent that handles code validation, case testing, and error analysis. The framework’s effectiveness is further enhanced by three specialized modules: an error handler that refines the debugging feedback, an example call detector that prevents runtime issues, and a function signature converter that provides a uniform template for the function head description. This structure enables PyCapsule to achieve high performance on widely used datasets, including HumanEval (Chen et al., 2021), HumanEval-ET (Dong et al., 2023), Mostly Basic Python Programming (MBPP) (Austin et al., 2021), MBPP-ET (Dong et al., 2023), and BigCodeBench  (Zhuo et al., 2024) with a high computational efficiency. The main contributions of our work are summarized as
  • Our PyCapsule is a resource-efficient framework that achieves high performance with significantly reduced computational overhead compared to existing multi-agent approaches. Our two-agent architecture demonstrates that complex coordination between multiple agents can be replaced by well-structured language prompts and modules without sacrificing accuracy.
  • Our efficient debugging mechanism aids the programmer agent through three specialized modules: an error handler, an example call detector, and a function signature converter. These achieve up to 5% accuracy improvement on HumanEval, 10% on HumanEval-ET using GPT models, and 25% on BigCodeBench using the Qwen 7B model compared to its 32B variant.
  • Through comprehensive analysis across multiple datasets and models, we observe a consistent decrease in the normalized success rate as we increase the number of self-debugging attempts. We hypothesize that this is caused by the increased problem complexity due to limited and noisy error messages for the self-debugging process.
2. Related Work
Automated high-quality code generation has evolved in several key directions related to prompt engineering (Knoth et al., 2024; Wei et al., 2022), multi-agent systems (Wooldridge and Jennings, 1995; Burgin and Dodig Crnkovic, 2009), and iterative debugging with runtime feedback mechanisms (Park et al., 2023; Chen et al., 2023; Islam et al., 2024; Huang et al., 2024b; Li et al., 2025). While LLM has demonstrated increasing capability in code generation, the code quality and correctness can be further enhanced by using intelligent debugging strategies. Therefore, previous works in self-debugging (Huang et al., 2023; Chen et al., 2023) show their capability in Automated Program Repair (APR) by identifying and fixing significant code flaws and bugs. Researchers have shown that ”large language models of code” (Huang et al., 2023) often struggle to directly fix bugs in the generated code due to the lack of task-specific fine-tuning. These fine-tuned models outperform the state-of-the-art APR tools (Xia and Zhang, 2022b; Fu et al., 2022; Joshi et al., 2023) but still suffer from the loss of pre-trained knowledge (Xia and Zhang, 2022a; Adnan et al., 2024), lack of debugging ingredients, inferior long-sequence handling ability, high resource constraints, long-tail problems, and multi-hunk error fixing (Huang et al., 2023).
Particulary, MapCoder (Islam et al., 2024) presents a multi-agent framework designed to emulate human-level programming processes. It employs four agents for retrieval, planning, coding, and debugging to perform iterative code generation and debugging. AgentCoder (Huang et al., 2024b) introduces a three-agent framework to address high token usage and coordination overhead observed in other multi-agent systems like MapCoder (Islam et al., 2024), MetaGPT (Hong et al., 2024), and ChatDev (Qian et al., 2024). The framework comprises a programmer agent, a test designer agent, and a test executor agent. The ”Debug Like a Human” framework (Zhong et al., 2024) introduces Large Language Model Debugger (LDB), which enhances the code quality by leveraging runtime execution information. LDB is performed iteratively through segmenting code, tracking intermediate variable states, and performing step-by-step verification with breakpoints. Moreover, some approaches have employed prompt engineering to improve the quality of code generation and debugging. For instance, CodeCoT (Huang et al., 2024a) leverages iterative refinement through Chain of Thought (CoT) (Wei et al., 2022) and logical reasoning. However, its capability is limited to refining syntax errors.
While these approaches have advanced automated code generations, they share several notable limitations, particularly the resource intensity for complex problems. Additionally, they often struggle to fully leverage error messages, which impairs their adaptability to nuanced debugging scenarios. Although this issue can be alleviated by generating complementary test cases, it can further increase resource consumption due to the reliance on preemptive test cases such that generation may become less reliable  (Li et al., 2025; Li and Yuan, 2024) and potentially require more complex agents.
To address these challenges, we introduce PyCapsule as a streamlined modular architecture with robust and optimized modules. It combines the knowledge from the operation research field and computer science to optimize the processing pipeline with merely two AI agents. This greatly reduces the computational overhead and facilitates more efficient and reliable code generation.
3. Methodology
3.1. Programmer and Executor Agents
Figure 1. PyCapsule framework as Python code generation service of OpenSI-CoSMIC (Adnan et al., 2024). PyCapsule consists of iterative code generation with debugging, error handling, and code execution within a Docker container.
PyCapsule employs two distinct types of agents as defined in the agent-based computing literature (Nwana, 1996; Jennings et al., 1998): smart agents and collaborative agents. The system consists of two agents: a programmer agent and an executor agent. The programmer agent generates code and handles debugging while the executor agent acts as an autonomous validator that continuously monitors and responds to code execution outcomes. When execution succeeds, it forwards the validated code to the user; otherwise, it generates detailed diagnostic reports and initiates a real-time feedback loop with the programmer agent for self-debugging.
The programmer agent leverages Ollama and GPT infrastructure embedded in OpenSI-CoSMIC (Adnan et al., 2024) to facilitate self-debugging and code generation through a structured pipeline. The programmer agent employs tailored system prompts with two programming modes, generation mode and fix mode. These prompts, detailed in Appendix  E, define the programmer agent’s persona (Shanahan et al., 2023), provide task-specific background, and structurally simplify code extraction. In the generation mode, the programmer agent uses CoT (Wei et al., 2022) to analyze the problem description and devise a structured solution to generate executable code. In the fix mode, the executor agent uses error information from the previous debugging attempt. This mode is supported by the original problem description, prior solution, and error messages processed by the error-handling module. The conversation history includes the most recent problem-solution pair. For MBPP with Qwen2.5-Coder-7B-Instruct, tests with 2 and 3 conversation pairs decrease the accuracy from 80.7% to 77.6% and 76.7%, respectively.
We implemented the executor agent using a lightweight Docker container. The container’s entry point is a shell script for installing required Python libraries and testing the generated code, which are generated by the programmer agent. Once the problem is solved, the container returns a success or failure status. The success status indicates that the generated code is correct by passing all the test cases while the failure status will trigger the error-handling module to refine and send the error message to the programmer agent. The programmer agent will then enter the fix mode to fix code flaws iteratively with self-debugging attempts. Particularly, with our analysis in Section 4.2 showing the diminishing influence of successive debugging attempts, the programmer agent is capable of refining the incorrect code with up to 5 self-debugging attempts. This seamless interaction between the two agents ensures robust and efficient feedback-driven debugging.
3.2. Supportive Modules
Existing approaches like MapCoder (Islam et al., 2024) and AgentCoder (Huang et al., 2024b) rely heavily on LLM agents for code generation. However, this requires a high resource consumption. In contrast, PyCapsule employs three specialised modules to handle each task deterministically, including a signature converter, an example call detector, and an error handling module.
The signature converter mitigates the instability of code generation by parsing problem descriptions provided in datasets through inferring structured function signatures. This module generates a function name, a robust signature, and an example call using the first test case for a given problem without revealing the case-testing result, thereby avoiding extra LLM calls for signature inference.
The example call detector ensures the safety of code execution by removing any internal example calls, which otherwise can lead to an infinite loop of function calls. This automatic detection replaces a dedicated LLM agent to check code execution safety.
The error-handling module refines error messages provided by the executor agent, giving concise and highly relevant natural language feedback. It also identifies the error type, filters irrelevant tracebacks, and resolves critical issues such as verbose recursion errors that can disrupt the programmer agent due to the LLM’s limited context length. This systematic approach requires fewer tokens in self-debugging iterations compared to the raw but redundant error messages.
These modules collectively enhance the code generation stability, safety, and self-debugging efficiency while reducing the computational overhead. More details are provided in Appendix C.
3.3. Code Generation Workflow
PyCapsule consists of three phases: code generation, execution, and self-debugging, shown in Figure 1. This pipeline integrates the programmer agent and executor agent with the self-bugging mechanism to ensure reliable code generation and case testing with high efficiency and accuracy.
The workflow starts with the programmer agent that receives a problem description, typically supplemented by a function signature either from the dataset or the signature converter module. Subsequently, the agent generates a response with CoT reasoning, environment setup instructions, and code synthesis. Then, the agent’s response is parsed to extract the synthesized code and environment setup instructions, with the latter automatically generating a text file, requirements.txt, containing all required Python libraries.
The extracted code is further processed through the example call detector to remove any unintended example calls. At the generation phase, the processed code with associated test cases from the dataset is saved in a Python entry file main.py for code execution. To ensure code execution safety, we add a timer with a maximum execution time to prevent infinite loops and raise an out-of-time error for those requiring longer running time.
At the execution phase, the executor agent runs main.py in a Docker container with Python libraries from requirements.txt installed. The corresponding code execution status, either successful or failed, determines whether the current generated code requires to be fixed by using the programmer agent’s fix mode. If so, error messages extracted from the executor agent are refined by the error-handling module and will used by the programmer agent to fix code flaws. This mode allows up to 5 self-debugging attempts, each incorporating code execution feedback from the previous bug-fixing process. The pseudo-code for the pipeline is provided in Appendix D.
4. Experiments
All GPU-related experiments were conducted on a single NVIDIA GeForce RTX 3090 (24 GB). We evaluated performance using success rate, defined by the ratio of correctly resolved problems to total problems. Our code will be released upon publication.
4.1. Dataset Overview
We used five benchmark datasets to evaluate PyCapsule: HumanEval (Chen et al., 2021), HumanEval-ET (Dong et al., 2023), MBPP (Austin et al., 2021), MBPP-ET (Dong et al., 2023), and BigCodeBench (Zhuo et al., 2024), where the ”ET” variants provide more test cases and enriched problem descriptions. These datasets encompass diverse programming challenges.
Particularly, HumanEval (Chen et al., 2021) consists of 164 human-curated programming problems for string manipulation, arithmetic, and basic data structures. Each problem includes a natural language description, a function signature, and test cases. MBPP (Austin et al., 2021) includes 974 programming problems spanning algorithms, data structures, and general-purpose tasks. It provides natural language descriptions and test cases but no function signatures to guarantee code generation stability. BigCodeBench (Zhuo et al., 2024) includes 1,140 problems with two splits: the complete split for code completion with specific docstrings and the instruct split with concise instructions for advanced reasoning.
4.2. Results and Analysis
Table 1. Success rate comparison on three popular LLMs. We provide the mean value and standard deviation for PyCapsule with three experiment repeats. The results of other methods are from their official reports.
Method HumanEval HumanEval-ET MBPP MBPP-ET BigCodeBench
GPT-4-Preview-1106
Direct       80.1  (Shinn et al., 2023)       81.1  (Islam et al., 2024)
CoT  (Islam et al., 2024) 89.0 82.4
Self-Planning (Islam et al., 2024) 85.4 75.8
AgentCoder [^1] Could not be reproduced during our experiments.  (Huang et al., 2024b) 96.3 86.0 91.8 91.8
MapCoder  (Islam et al., 2024) 93.9 82.9 82.1 57.7
PyCapsule (ours)       96.5$\pm$0.7       96.3$\pm$1.2       88.2$\pm$0.3       73.0$\pm$0.8
GPT-3.5-Turbo-1106
Direct       48.1  (OpenAI et al., 2024)       49.8  (Islam et al., 2024)
Direct (LDB (Zhong et al., 2024)) 73.8 67.6
CoT (Islam et al., 2024) 68.9 54.5
Self-Planning (Shinn et al., 2023) 60.3 55.7
Reflexion (Shinn et al., 2023) 67.1 73.0
AgentCoder [^1] Could not be reproduced during our experiments.  (Huang et al., 2024b) 79.9 77.4 89.9 89.1
MapCoder  (Islam et al., 2024) 80.5 70.1 78.3 54.4
LDB  (Zhong et al., 2024) 82.9 76.0
PyCapsule (ours)       85.2$\pm$0.7       84.7$\pm$0.6       78.5$\pm$1.2       62.4$\pm$0.6
Qwen2.5-Coder-7B-Instruct
Direct  (Hui et al., 2024) 88.4 83.5 41.0
PyCapsule (ours)       94.1$\pm$1.3       93.3$\pm$0.6       80.7$\pm$0.9       63.6$\pm$0.6       65.4$\pm$0.8
Unlike other multi-agent architectures discussed in Section 2 that heavily rely on code generation agents, PyCapsule has more effective coordination among the computational overhead reduction, LLM’s API calls, and token processing.
In Table 1, PyCapsule achieves the state-of-the-art code generation tasks across HumanEval, HumanEval-ET, and BigCodeBench. While showing strong performance on MBPP, it falls short in comparison to AgentCoder and Qwen-2.5 (Hui et al., 2024) which employs three-shot prompt engineering to stabilize code generation. Although it shows remarkable potential, we maintain a consistent prompting approach across datasets to ensure comparability and utilise our signature converter module to generate actionable function signatures. Remarkably, on HumanEval, the integration of Qwen-2.5-Coder-Instruct’s 7B model with PyCapsule achieves a 94.1% success rate compared to 92.7% by its 32B model. Furthermore, on BigCodeBench, PyCapsule using Qwen-2.5-Coder-Instruct’s 7B shows a 25% improvement over the 49.6% reported by the 32B model variant. This demonstrates a significant savings in computational power while exceeding the performance.
Figure 2. Distribution of relative success ratios along the self-debugging attempts. The relative success ratio refers to the number of successful test cases with given attempts over the total number of successful test cases, measured by unit %. Results on all the datasets consistently show a decreasing relative success ratio as the number of attempts increases.
Influence of Self-debugging Attempts.
To evaluate the influence of each self-debugging attempt on the success rate, we analyze the normalized and accumulated impacts of each incremental attempt using GPT-4-1106-preview, GPT-3.5-Turbo-1106, and Qwen-2.5-Coder-Instruct models as the LLM for code generation. The success rate normalization was performed as follows: given $N$ problems, the number of successfully solved problems is denoted as $S_{0}$ while the rest $N_{1}=N-S_{0}$ unsolved problems will be further resolved with a self-debugging attempt. Accordingly, the number of problems solved with $i\geq 1$ attempts is denoted as $S_{i}$, and $N_{i+1}=N_{i}-S_{i}$ for the rest. The independent influence is defined as
$I_{i}=\frac{S_{i}}{N_{i}},\quad\text{where }N_{i}=N-\sum_{j=0}^{i-1}S_{j}\ .$
Figure 3. The left subfigure shows the normalized influence of each self-debugging attempt through independent accuracy improvements. The right subfigure highlights the mean influence of each self-debugging attempt on the overall accuracy through the accumulated influence without normalization.
This metric highlights the independent contribution of each attempt. In Figure 3, the descending trend in the independent influence indicates diminishing effectiveness of debugging attempts, where problems requiring more attempts may need enriched conversation history or enhanced prompts to improve the accuracy. We hypothesize that the first few self-debugging attempts address the most significant and easily identifiable errors while substantially improving the success rate. However, since these primary issues are resolved, more subtle and complex issues occur, and they are difficult to be detected and solved even with more attempts, resulting in diminishing normalised accuracy while increasing the overall accuracy by small amounts, as demonstrated by the right plot of Figure 3 . Additionally, the accumulation of minor fixes may introduce unseen issues that further degrade the effect of self-debugging attempts.
4.3. Discussion
PyCapsule’s single-attempt accuracy is significantly lower than Qwen-2.5’s reported performance, and it varies considerably across datasets, as shown in Figure 2 and detailed in Appendix A. Nevertheless, PyCapsule achieves higher single-attempt accuracy compared to GPT models implemented in MapCoder and AgentCoder. This nuanced performance difference highlights the varying impacts of prompt engineering and framework structure on code generation without self-debugging. With more debugging attempts, PyCapsule shows overall remarkable improvements, surpassing all the other self-debugging methods. This iterative improvement highlights the effectiveness of prompt engineering and error-handling ability of PyCapsule. Prior works (Wei et al., 2022; Chen et al., 2024; Patel et al., 2024) have demonstrated how prompt design significantly influences model performance, especially for generative AI tasks. PyCapsule leverages sophisticated system prompts that are dynamically augmented in the code generation and fix modes. Furthermore, an important advantage of PyCapsule is its reduced reliance on LLM-based agents, leading to reduced token consumption and improved computational efficiency. Existing frameworks like MapCoder can make up to 17 API calls with GPT models per problem on HumanEval and 12 calls on MBPP, following $k\times t$ debugging attempts, where $k$ is the number of retrieved exemplars and $t$ is the number of self-debugging attempts. In contrast, PyCapsule streamlines this process by limiting the programmer agent with a maximum of 5 attempts, requiring at most 6 LLM API calls for each problem because planning and code generation are handled within a single API call. Our empirical analysis shows that PyCapsule achieves superior code generation ability while maintaining efficient API usage, requiring an average of only 2.09, 1.38, and 1.55 API calls for HumanEval using GPT-3.5-Turbo-1106, GPT-4-Preview-1106, and Qwen-2.5-Coder-Instruct-7B, respectively. This optimization demonstrates the possibility of achieving high performance with significantly reduced computational overhead.
5. Conclusion
We propose PyCapsule for significantly advancing reliable and automated code generation by integrating two AI agents with high-efficiency self-debugging modules. Its two-agent architecture effectively balances code generation automation and control while reducing resource consumption. Experimental results show that PyCapsule achieves high performance across multiple datasets, surpassing existing debugging-based approaches. Notably, PyCapsule enhances accuracy while using smaller language models and fewer API calls, offering a cost-effective alternative without compromising its performance. The iterative self-debugging feedback and dynamic prompt augmentation contribute to its robustness and improvements.
Appendix
Appendix A Contribution of Each Debugging Attempt
To further illustrate the influence of iterative debugging, we present a detailed breakdown of the success rate at each debugging attempt. Table 2 quantifies the proportion of problems successfully resolved at each stage. While the independent influence of successive attempts gradually decreases, the cumulative effect leads to an overall increase in the accuracy of problem-solving. Each debugging attempt contributes to solving additional problems, underscoring the value of iterative refinement in improving outcomes. This dual perspective—rising cumulative accuracy alongside diminishing returns per attempt—demonstrates the significance of leveraging efficient debugging strategies, such as enriched conversation history and refined prompts, to maximize the potential of LLM-based code generation systems.
Table 2. Accuracy of PyCapsule framework across code generation and debugging attempts (0–5) for OpenAI GPT-4-Preview-1106, GPT-3.5-Turbo-1106 and Qwen 2.5 Coder Instruct. Attempt 0 represents the initial solution generated by the Programmer agent, while subsequent attempts (1–5) indicate iterative fixes during the debugging process. The table reports accuracy with associated sample standard deviation (±$\pm s$). The results also showcase the incremental improvement achieved through PyCapsule’s iterative feedback mechanism.
Attempt HumanEval HumanEval-ET MBPP MBPP-ET BigCodeBench
GPT-4-Preview-1106
0 92.0 $\pm$ 1.4 88.0 $\pm$ 3.7 70.6 $\pm$ 1.0 67.5 $\pm$ 0.8 -
1 3.8 $\pm$ 0.7 8.2 $\pm$ 4.3 18.0 $\pm$ 1.1 17.2 $\pm$ 1.6 -
2 1.9 $\pm$ 1.1 2.3 $\pm$ 1.4 5.9 $\pm$ 1.5 8.6 $\pm$ 0.5 -
3 1.1 $\pm$ 0.4 0.8 $\pm$ 0.4 3.0 $\pm$ 0.3 3.2 $\pm$ 0.3 -
4 1.1 $\pm$ 0.7 0.6 $\pm$ 0.0 1.4 $\pm$ 0.5 1.9 $\pm$ 0.2 -
5 0.2 $\pm$ 0.4 0.0 $\pm$ 0.0 1.2 $\pm$ 0.4 1.6 $\pm$ 0.2 -
GPT-3.5-Turbo-1106
0 81.9 $\pm$ 0.6 77.2 $\pm$ 1.5 74.6 $\pm$ 1.8 70.3 $\pm$ 1.3 -
1 9.7 $\pm$ 0.3 13.2 $\pm$ 0.4 15.1 $\pm$ 0.9 16.2 $\pm$ 0.2 -
2 4.0 $\pm$ 0.4 5.3 $\pm$ 3.0 5.2 $\pm$ 0.7 6.6 $\pm$ 1.5 -
3 1.2 $\pm$ 0.8 1.9 $\pm$ 1.1 2.6 $\pm$ 0.5 3.3 $\pm$ 0.1 -
4 2.1 $\pm$ 0.0 1.4 $\pm$ 0.7 1.3 $\pm$ 0.4 2.0 $\pm$ 0.1 -
5 1.0 $\pm$ 0.4 1.0 $\pm$ 1.1 1.3 $\pm$ 0.8 1.5 $\pm$ 0.2 -
Qwen2.5-Coder-7B-Instruct
0 79.3 $\pm$ 1.4 81.3 $\pm$ 0.4 60.9 $\pm$ 0.7 59.7 $\pm$ 1.2 67.2 $\pm$ 0.9
1 14.9 $\pm$ 2.5 13.1 $\pm$ 2.0 22.4 $\pm$ 0.4 21.6 $\pm$ 0.7 14.9 $\pm$ 0.9
2 2.6 $\pm$ 0.7 2.2 $\pm$ 0.8 7.5 $\pm$ 0.2 8.7 $\pm$ 0.4 7.1 $\pm$ 0.5
3 1.5 $\pm$ 0.4 1.7 $\pm$ 1.0 4.5 $\pm$ 0.5 5.0 $\pm$ 0.3 4.7 $\pm$ 0.8
4 0.6 $\pm$ 0.6 1.5 $\pm$ 0.7 2.9 $\pm$ 0.3 2.9 $\pm$ 0.3 3.4 $\pm$ 0.7
5 1.1 $\pm$ 1.0 0.2 $\pm$ 0.4 1.8 $\pm$ 0.3 2.2 $\pm$ 0.5 2.7 $\pm$ 0.2
Appendix B HumanEval Experiments
Figure 4 and Figure 5 provide a detailed analysis of debugging effectiveness across multiple attempts in the HumanEval benchmark. By tracking the resolution process of individual tasks, they highlight the impact of iterative debugging in improving accuracy.
Figure 4. Visualisation of Debugging Attempts. The left column represents the number of debugging attempts on the HumanEval benchmark. Each point represents a problem, with failed tasks highlighted in red. The right column displays the mean number of debugging attempts and the associated sample standard deviations across three experiments for each model. The shaded areas in the right column indicate the variability in performance, illustrating the consistency and reliability of each model.
Figure 5. HumanEval experiment results using GPT-4-1106 across three repeats. Each line represents the progression of debugging attempts (up to 5) for tasks in the HumanEval dataset. Red markers denoting failed attempts. The coloured lines correspond to Experiment 1 (blue), Experiment 2 (green), and Experiment 3 (orange). The y-axis reflects the number of attempts taken to resolve each task.
Appendix C Modules (Detailed)
C.1. Signature Converter
The signature converter addresses code generation instability due to the lack of structured function signatures, particularly in the MBPP dataset (discussed in Section  4). Using the first test case, it infers the function name, input arguments, and constructs a robust function signature without revealing the expected output of the test case. The generated output includes:
  • The inferred function name.
  • A structured function signature (e.g., argument names and types).
  • An example function call derived from the test case.
For instance, given a test case like assert foo(4) == 16, the module generates:
### Required function name for your reference ‘foo()’
### Function signature for your reference - foo(arg_int: int)
### An example function call from private test cases - foo(4)
While return types are excluded from the function signature, they can be inferred from the test case’s data type. This module enhances the clarity and consistency of code generation.
C.2. Example Call Detection
The example call detection module identifies and removes instances where the programmer agent includes sample function invocations within its generated code output. Despite explicit system prompts to avoid such calls, they frequently appear in outputs, as observed in 21 of 164 HumanEval problems and 44 of 344 MBPP problems in one of our experiments with GPT-3.5-Turbo-1106. These embedded example calls pose a security risk as they can execute automatically during test case evaluation, potentially bypassing the runtime safety mechanisms we’ve implemented for controlled code execution.
This module scans the generated solution and removes any detected example calls before execution, ensuring system reliability. For example, calls like - foo(4) embedded in code are automatically removed. This safeguard prevents unintentional execution of potentially unsafe code outside the controlled environment.
C.3. Error-Handling
The error-handling module refines error messages returned by the executor agent, making them concise and supplemented with natural language feedback. It mimics human debugging practices by streamlining error messages into relevant components. The module processes error messages in the following steps:
  • Error Type Analysis: First, the module identifies error types and provides contextual feedback. For example, in the case of an AssertionError, the message might include: ”Your generated solution failed a test case. Please improve the logic of your solution.”
  • Relevance Filtering: It shortens error messages to focus on errors within the main.py file, truncating irrelevant tracebacks referencing external files.
  • Critical error-handling: For issues like RecursionError, the module truncates verbose error messages (e.g., maximum recursion depth exceeded) to ensure compatibility with the LLM’s context length. Without this, such errors could disrupt the LLM’s input pipeline.
This module improves the debugging process by filtering and contextualizing errors, ensuring that the Programmer Agent receives actionable and relevant feedback.
Appendix D Pseudo-code of PyCapsule
Algorithm 1 Pseudo-code of PyCapsule
Input: A user-defined query $\mathcal{Q}$ for code generation with a problem description, test cases, and the maximum number of self-debugging attempts $N=5$.
Output: A generated code satisfying the request in the problem description if all test cases passed, otherwise a failed status message.
Procedure:
1:Set and initialize a self-debugging attempt counter $i=0$.
2:Set an initial system prompt for generating code, a buffer $\mathcal{B}$ for storing conversation history, and a user prompt initialized by $\mathcal{Q}$.
3:Use $\mathcal{Q}$ and the LLM in PyCapsule to generate the function signature required by case testing.
4:Step 1: Generate code (”Generation Mode”).
5:Generate a prompt by integrating the system prompt and user prompt.
6:Use the prompt and the same LLM to generate a response for code generation.
7:Extract the required code, that is the function implementation, from the response.
8:Create a Python container with the required packages installed. If it exists, skip this step.
9:Test all the cases in the container using the extract code, and return the code execute status.
10:if all the test cases passed then
11:     Return the extracted code as the optimal generated code.
12:else if $i>N$ then
13:     Return the failed message extracted from the code execute status.
14:else
15:     Step 2: Code Debugging (”Fix Mode”).
16:     Extract the error message from the code execute status and add it to $\mathcal{B}$.
17:     Generate an error-handling prompt using the error-handling module.
18:     Update the system prompt for code debugging.
19:     Update the user prompt by integrating the error-handling prompt and messages in $\mathcal{B}$.
20:     Update $n=n+1$.
21:     Repeat Step 1.
22:end if
Appendix E System Prompts
E.1. Code Generation Mode
You are an experienced Python developer. Your task is to complete the function based on the provided function signature and description
- Analyze the description and provide a step by step reasoning on how to solve the problem.
- Maintain the function signature: Do not alter/modify the function signature.
- Avoid example calls: Do not include any example call in your response.
- If external libraries are needed, add a ’### Requirements Section listing them separated by ’,’. e.g. pandas, pyhton-dotenv ). If no external libraries are needed, add None’.
- Import all required libraries and complete the function in a ’### Code Section, enclosed with triple backticks.
Example Response Structure-
### Step-by-step reasoning
\$reasoning
### Requirements
\$external_libraries
### Code
\$code
E.2. Fix Mode
You are an experienced Python developer.
- Your previous solution resulted an error.
- Error message from a python compiler and Conversation history has been added for your reference.
- Analyze the description and provide a step by step reasoning on how to solve the problem.
- Maintain the function signature: Do not alter/modify the function signature.
- Avoid example calls: Do not include any example call in your response.
- If external libraries are needed, add a ’### Requirements Section listing them separated by ’,’. e.g. pandas, pyhton-dotenv ). If no external libraries are needed, add None’.
- Import all required libraries and complete the function in a ’### Code Section, enclosed with triple backticks.
Example Response Structure-
### Step-by-step reasoning
\$reasoning
### Requirements
\$external_libraries
### Code
\$code